A store wants to construct rectangular parking lot on land bordered on one side by a highway. It has 280 ft of fencing that is to be used to fence off the other three sides. What should be the dimensions of the lot if the enclosed area is to be a maximum?
Answer by checkley77(12569)

THE MAXIMUM AREA OF A 3 SIDED LOT IS 1:2:1
THE SUM OF THESE RATIOS IS 4
DIVIDE 280/4=70
THUS THE ANSER IS 70:140:70
PROOF
70*140=9,800 FT^2
69*142=9,798 FT^2
71*138=9,798 FT^2